Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labor-h13. What is the spending variance related to advertising? (Indicate the effect of each variance by selecting F for favorabl14. What is the spending variance related to sales salaries and commissions? (Indicate the effect of each variance by selecti15. What is the spending variance related to shipping expenses? (Indicate the effect of each variance by selecting F for fa
Answer #1

13) Spending variance in Advertising = 220000-235000 = 15000 U

14) Spending variance in sales salaries and commission = (24600*14+140000)-465000 = 19400 F

15) Spending variance in Shipping expense = (24600*5-135000) = 12000 U